Miyo, YamaT, Uchiyama, Akutsu
Cont'd from 16126. Like to replace the oplev TX holders.
- For MCi, I confirmed the oplev light path was not clipped, and fixed the TX and also RX, which were not fixed yet; note that TX hoder replacing had been done already.
- For MCo, I replaced the TX holder, and align the oplev light beam, and fixed the TX holder. Note that before this work, the MCo oplev has been already centered (zero-ed). See 16225.
- Finally, in order to remember today's alignment of MCi and MCo, we tweaked the TX side of MCo so that the corresponding oplev output got zero. So do MCi. Note that for the case of MCi, whichever of TX aand RX are ok to be tweaked, as I had already broken the reference.